r its Arizona facility. The machine costs $158,000 and is expected to have a useful life of 9 years, with a terminal disposal value of $3,000. Savings in cash operating costs are expected to be $37,500 per year. However, additional working capital is needed to keep the machine running efficiently. The working capital must continually be replaced, so an investment of Lab's required rate of return is 12%. Ignore income taxes in your analysis. Dunkin Lab plans to purchase a new centrifuge machine Present Value of $1 table Present Value of Annuity of $1 table Future Value of $1 table Future Value of Annuity of $1 table
Calculate net present value. 1. Calculate internal rate of return 2. 3. Calculate accrual accounting rate of return based on net initial investment. 4. Calculate accrual accounting rate of return based on average investment. 5. You have the authority to make the purchase decision. Why might you be reluctant to base your decision on the DCF methods?
Answer #1

Initial Investment = $158000 + $12000 = $170000
Savings = $37500
Depreciation = ($158000 - $3000) / 9 = $17222
Net Income = $37500 - $17222 = $20278

1. PV annuity @12% for 9 years = 5.3282, PV factor for 9th year @12% = 0.3606
NPV = $37500 x 5.3282 + $3000 x 0.3606 + $12000 x 0.3606 - $170000 = $35216.50

2. For IRR, NPV = 0,
PV annuity @17% for 9 years = 4.4506, PV factor for 9th year @12% = 0.2434
NPV = $37500 x 4.4506 + $3000 x 0.2434 + $12000 x 0.2434 - $170000 = $548.50

PV annuity @18% for 9 years = 4.3030, PV factor for 9th year @12% = 0.2255
NPV = $37500 x 4.3030 + $3000 x 0.2255 + $12000 x 0.2255 - $170000 = ($5255)

RR = 17% + ($548.50) / ($548.50+$5255) = 17.09%

3. Accounting Rate of Return = $20278 / $170000 x 100 = 11.93%

4.

Year Book Value
0 170000
1 152778
2 135556
3 118334
4 101112
5 83890
6 66668
7 49446
8 32224
9 2
910010


Accounting Return = $20278 / $91000 x 100 = 22.28%
Average Investment = $910010 / 10 = $91000

5. Since the project gives positive NPV and IRR higher than required rate of return, project is accepted.